OpenStudy (anonymous):

Solve using substitution. y = 2x + 3 5x - y = -3 can someone show work?

Parth (parthkohli):

Which means,\[\rm 5x - (2x + 3) = -3\]Right?

OpenStudy (anonymous):

I have no idea, i get confused

Parth (parthkohli):

\[\rm 5x - 2x - 3 = -3\]Right?

OpenStudy (anonymous):

Since \[y=2x+3\] , you have to plug this into \[5x-y=-3\]..